If you are being supported by your parents, use the total amount of support you are getting as your income. If you are living with your parents, try to estimate Your portion of the household expenses. For example, if you use one room of their 5-room house, find out the monthly mortgage payment that your parents make and divide that number by five to get the rent expense for the room that you are using (This will be your rent amount). Use similar logic for all the other expenses that they are paying while you are living at home such (utilities) the light, gas, and water bill; insurance for car and groceries and etc. I want you to (experience) get an idea of how much it costs for you to live your life for a month.

There are four sections for this assignment:

1. Following the example provided in class: Prepare a Budget (what you think you will earn and spend) for the month of February 2021. After you make the budget, keep track of your actual expenses throughout the entire month. Then after the month is over, you will compare the Actual against your Budget to create Variances.

Make sure the Budget has three columns (Budget, Actual, and Variance). Keep track of your income and expenses. You will need these actual numbers to fill in the “Actual” column to calculate your variances. (5 points)

2. Following the examples provided in class: Prepare your Personal Income Statement for the month ending February 28th, 2021. (10 points)


3. Prepare your Personal Balance Sheet as of February 1st, 2021. (10 points)

4. Using your financial statements Calculate the following ratios: (10 points)

a. Current ratio

b. Month’s living expenses covered ratio (Emergency fund ratio)

c. Debt ratio

d. Inverse debt coverage ratio (look it up in the book)

e. Savings ratio

f. Tell me what you think (interpret) about each of your ratios and what can you do to improve them (if necessary)?


1
Expert's answer
2021-01-28T04:00:38-0500



The planned data is the budget, actual data - actual. The deviation of the fact from the plan is Variance.

"Variance=budget- actual"

The balance is the difference between income and expenses.

a. Current ratio

"Current Ratio= \\frac{Current assets}{Current liabilities}=\\frac{20000}{19033.33}=1.0507\n\u200b\t \n\u200b"

b. Month’s living expenses covered ratio (Emergency fund ratio)

"Months’ living expenses covered ratio =\frac {total liquid assets}{monthly total cash outflows}==\frac{20000}{19033.33}=1.0507"

c. Debt ratio

"Debt ratio=\\frac{Total debt}{Total assets}=\\frac{7000+833.33+5200}{20000}=0.68\n\n\u200b\t \n\u200b"

d. Inverse debt coverage ratio

"Debt Coverage Ratio=\frac{CFO}{Total Debt}=\frac{966.67}{7000+833.33+5200}=0.07"

e. Savings ratio

"Savings ratio=\frac{balance}{income}=\frac{966.67}{20000}=0.048"

f. judging by the calculated coefficients, the income covers the expenses, but the balance is small, so it is necessary to either reduce the income or reduce the expenses